OpenStudy (anonymous):

HOW MUCH IS EACH AT REGULAR PRICE? NOT INCLUDING TAX. James bought two T-shirts and one pair of jeans at an online store and paid $40, not including taxes, for his purchase. A month later, the same store sold the T-shirts and jeans at a 50% discount from their original prices. James bought two T-shirts and five pairs of jeans for $60, not including taxes.

OpenStudy (anonymous):

J = $20 for a pair of jeans

OpenStudy (anonymous):

T= the price of a t shirt, J= the price of a pair of jeans (1) 2T + J = 40 Now a month later the prices for the t shirt and jeans were 50% off which is half the price so: 2(T/2) + 5(J/2) = 60, which can be written as: (2) T + (5/2)J = 60 Now we have two equations. Take equation (1) and solve for J: J = 40 - 2T, which you can plug into equation (2) to get: T + (5/2)(40 - 2T) = 60, which if you distribute and combine like terms: T + (200/2) - 5T = 60 4T = 40 T = $10 for a t shirt, now we can plug into equation (1) 2(10) + J = 40, solve for J and you get J = $20 for a pair of jeans

OpenStudy (anonymous):

Thanks!!